Figure 1: L^{A}TEX single-column page layout. The actual proportions correspond to parameter values in the 11pt BOOK document style. Note that standardconforming DVI drivers are required to place the TEX upper-left page corner one inch over and down from the corner of the physical output page. This figure is scaled to 50% of actual page size. It was produced on March 25, 2011 at 15:43.

 [figure: 1in; 1in; A line of text. . .; Next line of text. . .; \baselinestretch×\baselineskip; \oddsidemargin; \evensidemargin; Page Text; Page Footer; \footheight; \footskip; \textwidth; \textheight→; \headheight; \headsep; \topmargin; Page Header; Margin note A; \marginparpush; Margin note B; \marginparwidth; \marginparsep]

Figure 2: L^{A}TEX double-column page layout. The actual proportions correspond to parameter values in the 11pt BOOK document style. Note that standardconforming DVI drivers are required to place the TEX upper-left page corner one inch over and down from the corner of the physical output page. This figure is scaled to 50% of actual page size. It was produced on March 25, 2011 at 15:43.

 [figure: 1in; 1in; A line of text. . .; Next line of text. . .; \oddsidemargin; \evensidemargin; ← rule width is \columnseprule; \baselinestretch×\baselineskip; \columnsep; Page Footer; \footheight; \footskip; \textwidth; \textheight→; \headheight; \headsep; \topmargin; Page Header; Margin note A; \marginparpush; Margin note B; \marginparwidth; \marginparsep]

Figure 3: Page layout parameters (in centipoints) for preceding figures. It was produced on March 25, 2011 at 15:43.

\BASELINESKIP  2000
\COLUMNSEPRULE  0
\COLUMNSEP  1000
\COLUMNTWOX  20374
\COLUMNXMIDDLE  19874
\DVIXOFFSET  7227
\DVIYOFFSET  7227
\EVENSIDEMARGIN  10841
\FOOTHEIGHT  2400
\FOOTNOTESEP  665
\FOOTSKIP  4207
\HEADHEIGHT  2400
\HEADSEP  1807
\HEADYORIGIN  56414
\INNERHEIGHT  64234
\INNERWIDTH  46976
\MARGINNOTEHEIGHT  4000
\MARGINNOTEXORIGIN 38935
\MARGINNOTEYA  42007
\MARGINNOTEYB  36007
\MARGINPARPUSH  2000
\MARGINPARSEP  2800
\MARGINPARWIDTH  7227
\ODDSIDEMARGIN  3613
\PAGEHEIGHT  79497
\PAGEWIDTH  61430
\PAGEXORIGIN  7227
\PAGEYORIGIN  8036
\PAGEYTOP  72270
\RIGHTEDGE  36135
\TAD  800
\TEXTHEIGHT  50400
\TEXTWIDTH  32522
\TEXTWIDTHY  16800
\TOPMARGIN  5420